OpenStudy (anonymous):

A company rents riding equipment for a fixed amount plus a fee based on the number of days for which the equipment is rented. The table shows the total charges, y, in dollars, of renting riding equipment for x number of days: Riding Equipment Rentals Number of days (x) Total Charges in dollars (y) 0 6 1 16 2 26 What is the fixed amount charged?

OpenStudy (anonymous):

can someone help

Vocaloid (vocaloid):

the fixed amount is the value of y when x = 0

OpenStudy (anonymous):

so would the answer be 6

Vocaloid (vocaloid):

yes

OpenStudy (anonymous):

can you help me with another one

Vocaloid (vocaloid):

sure

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Barney wants to rent a tent. He has to pay a fixed base cost plus a daily rate for renting the tent. The table shows the amount of money, y, in dollars, that Barney has to pay for renting the tent for x days: Tent Rental Number of days (x) Rent (dollars) (y) 0 3 1 8 2 13 3 18 4 23 Which equation best shows the relationship between x and y? y = 5x + 3 y = x + 8 y = 5x + 8 y = x + 3

Vocaloid (vocaloid):

your equation will be in the form y = mx + b let's find m first. when x increases by 1, what value does y increase by?

OpenStudy (anonymous):

i don't know what m would be

Vocaloid (vocaloid):

do you see the pattern in the y values in the table?

OpenStudy (anonymous):

i can find the pattern

Vocaloid (vocaloid):

every y value increases by a certain number each time, what is that number?

OpenStudy (anonymous):

it goes up by 5

Vocaloid (vocaloid):

great, so that means our m value (slope) is 5 so right now we have y = 5x + b, now let's find b b is the y value when x = 0, so what is b?

OpenStudy (anonymous):

would it be 1/5

Vocaloid (vocaloid):

m = 5, since y increases by 5 every time x increases by 1 right now, we have y = 5x + b, now we need to find b b is the y value when x = 0, so what is b?

OpenStudy (anonymous):

i don't know what be would be

Vocaloid (vocaloid):

what is the value of y when x = 0?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

ohh yea i forgot that b was the y intersect my bad

Vocaloid (vocaloid):

right, so b = 3 putting everything together: y = mx + b = 5x + 3 and that's our answer

OpenStudy (anonymous):

ok thank you sooooooooo much